What will be the value of x after the following code runs:

func (y) {

while y > 0

  y--

  x += y

}

x = 0;

func (5)
Answer #1

Ans) 4

Explanation:

Here y =5 as function parameter .

The condition inside While evaluates true first statement inside While is y-- i.e 5-1 =4 and x=0 finally statement is x+=y

I.e x=x+y

X=0+4=4

Answer returns as 4